# 3D-Properties: Identifying Challenges in DPO and Charting a Path Forward

### v1

@@ D.1 Training Setting
In this section, we provide a detailed overview of our training settings. Following the implementation of [23], we use the Adam optimizer with the learning rate set to 5e-7, and the default value for $\beta$ was set to 0.1. We set the batch size to 80 and the number of gradient accumulation steps to 2. The training epoch was set to 1. In IPO training, we set $\eta$ to be 0.1. In SLiC training, we set $\delta=5$, $\eta=0.1$. All experiments were conducted on a cluster consisting of 40 A100 GPUs.
@@ D.2 Supplementary Experimental Results
Table 8 and Table 9 show the performance enhancement in MATH and SuperCLUE by vanilla DPO training respectively. It is easy to see that Scenario 1 where both the chosen responses and the rejected responses are on-policy performs best.
Figure 9 and Table 10 represent the additional results on DPO variants and regularization techniques. It can be seen that DPO variants can achieve on-par or better performance compared with vanilla DPO. Figure 9 shows that as $\beta_{2}$ decreases, the performance in poem generation initially improves, reaches the peak point at around $\beta_{2}=0.08$, and subsequently declines. The initial improvement is intuitive. We conjecture that an excessively low $\beta_{2}$ may cause DPO to perform like SFT on the chosen responses, thereby reducing its generalization capabilities. For different tasks, the peak point of $\beta_{2}$ can be different. For example, in mathematical reasoning, we can set $\beta_{2}$ to be $0.01$ and achieve better performance than vanilla DPO.
@@
Table 8: Vanilla DPO: Baichuan2-13B accuracy on MATH and SuperCLUE.
MATH SuperCLUE
5 points 4&5 points 5 points 4&5 points
basemodel 6.0% 12.2% 46.3% 58.8%
Scenario 1 7.9% 14.4% 52.8% 64.6%
Scenario 2 4.8% 9.2% 47.9% 61.8%
Scenario 3 4.3% 12.8% 41.2% 57.0%
Scenario 4 3.2% 9.3% 39.5% 52.9%
Table 9: Vanilla DPO: Baichuan2-33B accuracy on MATH and SuperCLUE.
MATH SuperCLUE
5 points 4&5 points 5 points 4&5 points
basemodel 25.7% 36.5% 79.5% 84.4%
Scenario 1 29.9% 37.5% 80.2% 86.6%
Scenario 2 29.2% 36.6% 72.2% 79.0%
Scenario 3 28.2% 37.3% 74.8% 84.9%
Scenario 4 28.6% 38.1% 80.2% 85.8%
Figure 9: Performance on poem generation, $\beta_{2}$ varying with $\beta_{1}=0.1$.
Table 10: DPO and its variants/regularized version performance on mathematical reasoning. In Flex-DPO, $\beta_{1}=0.1$, $\beta_{2}=0.01$.
@@
Figure 11: Comparison between DPO and reward model training on the training set of MATH. As can be seen, in the second epoch of DPO training, the loss is very small and the accuracy of distinguishing the chosen response from the rejected response is 100%, which indicates that the model overfits the training data.
@@
Figure 12: Comparison between DPO and reward model training on the test set of HH-rlhf.

### v2

@@ D.1 Training Setting
In this section, we provide a detailed overview of our training settings. Following the implementation of Rafailov et al. (2024), we use the Adam optimizer with the learning rate set to 5e-7. The most sensitive parameter in the DPO algorithm is $\beta$ (and learning rate but less significant). Here we use the default setting aligned with the original DPO paper, The $\beta$ here is set to be 0.1 and the learning rate is set to be $5\times 10^{-6}$, which is the best hyperparameter set as far as we explored. We set the batch size to 80 and the number of gradient accumulation steps to 2. The training epoch was set to 1. In IPO training, we set $\eta$ to be 0.1. In SLiC training, we set $\delta=5$, $\eta=0.1$. All experiments were conducted on a cluster consisting of 40 A100 GPUs.
@@ D.2 Supplementary Experimental Results
Table 7 reports the log probabilities before and after training. Scenario 1 exhibits the slowest decline in likelihood compared to other scenarios, effectively mitigating the adverse effects of 3D-properties.
Table 8 show the performance enhancement in MATH and SuperCLUE by vanilla DPO training respectively. It is easy to see that Scenario 1 where both the chosen responses and the rejected responses are on-policy performs best.
Figure 4 and Table 9 represent the additional results on DPO variants and regularization techniques. It can be seen that DPO variants can achieve on-par or better performance compared with vanilla DPO. Figure 4 shows that as $\beta^{-}$ decreases, the performance in poem generation initially improves, reaches the peak point at around $\beta^{-}=0.08$, and subsequently declines. The initial improvement is intuitive. We conjecture that an excessively low $\beta^{-}$ may cause DPO to perform like SFT on the chosen responses, thereby reducing its generalization capabilities. For different tasks, the peak point of $\beta^{-}$ can be different. For example, in mathematical reasoning, we can set $\beta^{-}$ to be $0.01$ and achieve better performance than vanilla DPO.
@@
Among the four scenarios, Scenario 1, where both chosen and rejected responses are on-policy, ensures a more stable DPO training process and delivers the best testing performance, as shown in Table 1. As shown in Table 7, Scenario 1 exhibits the slowest decline in likelihood compared to other scenarios, effectively mitigating the adverse effects of 3D-properties. This explains the superior performance of on-policy DPO in our tests.
In addition to pure on-policy and off-policy DPO, Scenario 2 where the chosen response is off-policy and the rejected response is on-policy is also commonly seen in the industry. For example, in some math problems, researchers used to treat the accurate solution in the dataset to be the chosen response and the wrong answer generated by the LLM itself to be the rejected response, which we show is harmful. Scenario 3 is a mirror experiment to Scenario 2. These additional experimental results confirm that as long as the off-policy data is mixed into the training preference dataset, the performance of DPO will be weakened.
Table 7: Impact of on-policy training data on the results. $\log\pi(a^{+})$ and $\log\pi(a^{-})$ represent the average log probability per token for the chosen and rejected responses, respectively.
Baichuan2-13B Baichuan2-33B
$\log\pi(a^{+})$ $\log\pi(a^{-})$ $\log\pi(a^{+})$ $\log\pi(a^{-})$
basemodel -0.9181 -0.9393 -0.3603 -0.3634
DPO in Scenario 1 -0.9681 -0.9982 -0.3629 -0.3670
basemodel -1.6776 -0.9238 -1.4314 -0.3525
DPO in Scenario 2 -1.6265 -1.2293 -1.2734 -0.4254
basemodel -0.9461 -1.7110 -0.3460 -1.1204
DPO in Scenario 3 -0.8786 -1.8848 -0.3439 -1.4333
basemodel -1.7468 -1.7110 -1.2838 -1.1451
DPO in Scenario 4 -1.6617 -1.8265 -1.2273 -1.2234
Table 8: Vanilla DPO: Baichuan2-13B and Baichuan2-33B accuracy on MATH and SuperCLUE.
Baichuan2-13B Baichuan2-33B
MATH SuperCLUE MATH SuperCLUE
5 points 4&5 points 5 points 4&5 points 5 points 4&5 points 5 points 4&5 points
basemodel 6.0% 12.2% 46.3% 58.8% 25.7% 36.5% 79.5% 84.4%
Scenario 1 7.9% 14.4% 52.8% 64.6% 29.9% 37.5% 80.2% 86.6%
Scenario 2 4.8% 9.2% 47.9% 61.8% 29.2% 36.6% 72.2% 79.0%
Scenario 3 4.3% 12.8% 41.2% 57.0% 28.2% 37.3% 74.8% 84.9%
Scenario 4 3.2% 9.3% 39.5% 52.9% 28.6% 38.1% 80.2% 85.8%
Table 9: DPO and its variants/regularized version performance on mathematical reasoning. In Flex-DPO, $\beta^{+}=0.1$, $\beta^{-}=0.08$.
@@
(a) Accuracy of RM and DPO on HH-rlhf eval set over the training process.
(b) Accuracy of RM and DPO on MATH eval set over the training process.
Figure 7: Comparison of DPO and RM Training. RM training demonstrates greater stability, while DPO training shows significant fluctuations.
@@
Figure 8: Comparison between DPO and RM training on the training set of MATH. As can be seen, in the second epoch of DPO training, the loss is very small and the accuracy of distinguishing the chosen response from the rejected response is 100%, which indicates that the model overfits the training data.
@@
Figure 9: Comparison between DPO and RM training on the test set of HH-rlhf.
D.3 Limitations
Despite the insights provided in this study, several limitations remain. First, while our theoretical analysis and experiments highlight the 3D-properties as a key factor in DPO’s suboptimal performance, the complexity of real-world LLMs may involve additional factors that were not fully explored. Second, our experimental evaluation, though spanning diverse tasks such as mathematical reasoning and instruction following, is limited in scope and may not generalize across all LLM applications. Additionally, the regularization techniques proposed, while effective in our controlled settings, require further validation in larger-scale models and more diverse datasets. Lastly, although we contrasted DPO with RM-based alignment, our study does not exhaustively address other potential reward-free methods, leaving open questions for future exploration.